OpenStudy (anonymous):

find the slope of the line passing through the given points (1,3) and (2,5)

OpenStudy (kira_yamato):

Slope = (5 - 3) / (2 - 1) = 2

OpenStudy (anonymous):

delat y over delat x.... 2 / 1 = 2

OpenStudy (anonymous):

*delta

OpenStudy (anonymous):

how do i graph this?

OpenStudy (precal):

plot the 2 given points and draw a line through them.